Income Statement

Year

  2019     

     2020

Net Revenue

140,000

- Cost of Goods Sold

  70,000

- Depreciation Expense

    9,000

EBIT

  61,000

- Interest Expense

  10,500

Income Before Taxes

  50,500

Tax Expense

  10,605

Net Income

2019 Dividend

  39,895

    9,974

Balance Sheet

Year (end of)

2019

   2020

2019

2020

Assets

Liabilities

Current Assets

Current Liabilities

        Cash and Equivalents

  10,000

        Accounts Payable

21,000

        Accounts Receivable

  25,000

Long-term Debt

95,000

        Inventory

  12,000

Total Liabilities

116,000

Fixed Assets, Net

165,000

Stockholders' Equity

Total Assets

212,000

Common Stock

44,000

Retained Earnings

52,000

Total Stockholders Equity

96,000

Sales will grow by 10% in 2020. All costs, assets, and current liabilities vary directly with sales. Interest Exp., Common Stock, Tax Rate and Div. payout ratio are constant. L-T Debt=Plug number.

1A. Prepare a 2020 forecast.  What is the 2020 Dividend and Addition to Retained Earnings?

1B. If a bank will allow Atlantic to borrow 2.5 times prior year EBITDA, how much total Long-Term Debt would the bank allow in 2020?

1C. What is Atlantic’s Days Accounts Payables in 2020?  By how many Days would A/P need to increase to balance the Balance Sheet if Long-Term Debt = $70,000?
Answer #1
Income Statement
2019 2020
Net Revenue          140,000    154,000 (1.1*140000)
Cost of goods sold            70,000       77,000 (1.1*70000)
Depreciation expense              9,000         9,900 (9000*1.1)
EBIT            61,000       67,100
Interest expense            10,500       10,500
Income Before taxes            50,500       56,600
Tax expense(21%)            10,605       11,886
Net Income            39,895       44,714
Dividend              9,974       11,179
Payout ratio=9974/39895 0.250006 0.250006
Tax Rate=10605/50500 0.21 0.21
Addition to Retained Earning            29,921       33,535 (44714-11179)
BALANCE SHEET
2019 2020 2019 2020
Assets: Liabilities:
Current Assets Current Liabilities:
Cash & Cash Equivalent            10,000       11,000 (1.1*10000) Accounts Payable            21,000       23,100 (1.1*21000)
Accounts Receivable            25,000       27,500 (1.1*25000) Long Term debts            95,000       80,565 (233200-129535-23100)
Inventory            12,000       13,200 (1.1*12000) Total Liabilities          116,000    103,665
Fixed Assets(Net)          165,000    181,500 (1.1*165000) Stockholders Equity:
Total Assets          212,000    233,200 Common Stock            44,000       44,000
Retained Earnings            52,000       85,535 (52000+33535)
Total Stockholders Equity            96,000    129,535
Liabilities+Equity          212,000    233,200
1A 2020 Dividend            11,179
2020 addition to retained earnings            33,535
1B 2019 EBIT            61,000
Depreciation expense              9,000
2019 EBITDA            52,000
Bank will allow to borrow=2.5*52000          130,000
1C DAYS ACCOUNT PAYABLE IN 2020
Accounts Payable Turnover=Cost of Goods sold/Accounts Payable
Accounts Payable Turnover= 3.33333333 (77000/23100)
Days Accounts Payable=365/Accounts Payable Turnover
DAYS ACCOUNT PAYABLE IN 2020 109.5 Days (365/3.3333)
If Long Term debt =$70000